OpenStudy (anonymous):

the ratio in a certain factory is 3 to 4. there are 204 men. how many workers are there?

OpenStudy (anonymous):

The ratio of men and women in a certain factory is 3 to 4. There are 204 men. How many workers are there? 3 + 4= 7 204/7= 29.14 ?????? --------------------- It's 3 men to 4 women. 3 out of 7 are men. 204 = 3/7 of the total 68 = 1/7 of the total So from that what would be the total?

OpenStudy (anonymous):

@jayybee7

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

68 divided by 7?

OpenStudy (anonymous):

I don't know how else to explain it but the total is 476.